# Graph is a Natural Regularization: Revisiting Vector Quantization for Graph Representation Learning

### v1

@@ C. Experiments
Datasets. We use both homophilous and heterophilous graphs in our experiments. To evaluate codebook utilization, we use various datasets, including Cora, CiteSeer, PubMed, Amazon-Computer, Amazon-Photo, WikiCS, Amazon-Ratings, and Roman-Empire. To assess transferability, we use cross-task and cross-domain datasets. Specifically, we use Cora, PubMed, and WikiCS for node classification; WN18RR and FB15K237 for link prediction; and HIV and PCBA for graph classification. Finally, we evaluate serialization ability using the same datasets employed for codebook utilization. Detailed dataset statistics are summarized in Table 4.
@@
Baselines. We use different baselines for three parts of our main experiments.
@@
  • Codebook Utilization. We primarily adopt codebook mitigation strategies originally developed in the vision and language domains, including EMA (Łańcucki et al. 2020), affine parameters (Huh et al. 2023), codebook reset (Zeghidour et al. 2021), and pretrained encoders (Zhao et al. 2024b). We further include SimVQ (Zhu et al. 2024), which addresses the codebook collapse via one-single MLP layer over the latent basis vectors. Additionally, we compare with HQA-GAE (Zeng et al. 2025), a recent graph VQ model that applies a hierarchical VQ structure and an annealing strategy for codeword selection.
@@
  • Transferability. To evaluate the effectiveness of RGVQ in learning transferrable graph tokens, we integrate it into a graph foundation model, i.e., GFT, and compare it with vanilla Graph VQ and its variants with different mitigation strategies. Moreover, we include supervised GNNs, i.e., GCN and GAT, and graph self-supervised methods, i.e., DGI (Veličković et al. 2018), BGRL (Thakoor et al. 2021), and GraphMAE (Hou et al. 2022). The supervised GNNs are trained directly on each target dataset, while the self-supervised methods and all GFT variants are pretrained on the full set of datasets and then fine-tuned per target.
@@
  • Serialization.To further evaluate the effectiveness of RGVQ in serialization, we integrate it into the Graph Quantized Transformer (GQT) (Wang et al. 2025), where discrete tokens serve as the input sequence to a vanilla Transformer backbone. We follow the original sequence reconstruction method and compare the performance of RGVQ-enhanced GQT against the original GQT, supervised GNNs, and graph transformers, including GraphGPS (Rampášek et al. 2022), SGFormer (Wu et al. 2023), and Exphomer (Shirzad et al. 2023).
Implementation Details. We provide the hyperparameters and implementation details of our main experiments.
@@
  • Codebook Utilization. The implementation of all collapse mitigation strategies is the same with the empirical study (see Appendix A). For HQA-GAE, we use one-head VQ model and use the same hidden dimension and number of GNN layers as other methods, while all remaining hyperparameters follow the original paper. For RGVQ, we set the codebook size $K$ to 512, and sample 20 contrastive examples for both the positive and negative sets. The training weights for the link prediction, node feature reconstruction, contrastive regularization, commitment loss, and vocabulary loss are set to 0.01, 100, 1, 0.1, and 0.9, respectively. We set the temperature for the Gumbel-Softmax trick to 0.1.
@@
  • Transferability. We use RGVQ as a plugin within the pretraining pipeline of GFT. Specifically, we retain the link, node feature, and tree embedding reconstruction tasks and integrate RGVQ with respective weights of 100, 1, 0.01, and 10. For all other configurations, including downstream fine-tuning in GFT, we strictly follow the settings from the original paper (Wang et al. 2024).
@@
  • Serialization.We pretrain RGVQ using the same configuration as in the codebook utilization experiments for 200 epochs. We then load the pretrained checkpoints, including the codebook, GNN encoder, and VQ module weights. We follow the hyperparameter settings from the original GQT paper (Wang et al. 2025).
D. Complexity Analysis
@@
Cora PubMed Citeseer Photo Computer WikiCS Ratings Roman Questions
Graph VQ
94.47
$\pm$8.65
4.14
$\pm$1.03
60.09
$\pm$5.59
1.00
$\pm$0.00
1.00
$\pm$0.00
10.18
$\pm$2.13
13.29
$\pm$2.89
10.84
$\pm$3.48
20.78
$\pm$3.65
EMA
91.68
$\pm$9.17
5.12
$\pm$1.46
55.15
$\pm$6.73
1.00
$\pm$0.00
1.00
$\pm$0.00
11.27
$\pm$3.36
9.12
$\pm$2.33
6.20
$\pm$3.24
14.15
$\pm$3.51
AP
75.32
$\pm$6.28
126.55
$\pm$12.64
9.03
$\pm$2.16
54.95
$\pm$5.43
59.33
$\pm$8.55
83.55
$\pm$9.86
73.82
$\pm$8.14
118.46
$\pm$16.21
66.57
$\pm$8.27
Reset
65.79
$\pm$8.56
102.78
$\pm$15.78
85.19
$\pm$4.41
10.73
$\pm$1.98
17.18
$\pm$2.37
134.44
$\pm$7.35
130.83
$\pm$8.88
150.51
$\pm$11.15
141.98
$\pm$10.11
PT
60.57
$\pm$10.25
6.17
$\pm$1.12
138.98
$\pm$10.54
3.78
$\pm$1.37
2.94
$\pm$1.27
3.10
$\pm$1.31
37.65
$\pm$5.76
14.49
$\pm$2.52
58.99
$\pm$8.34
SimVQ
40.09
$\pm$6.53
23.96
$\pm$2.56
38.11
$\pm$6.67
37.29
$\pm$4.85
40.47
$\pm$6.54
45.90
$\pm$7.35
16.08
$\pm$4.11
42.22
$\pm$8.34
21.71
$\pm$5.27
HQA-GAE
130.06
$\pm$5.52
164.77
$\pm$14.15
93.67
$\pm$11.32
166.32
$\pm$10.98
114.08
$\pm$10.15
98.73
$\pm$7.82
92.17
$\pm$8.66
89.05
$\pm$8.23
72.86
$\pm$7.79
RGVQ
211.69
$\pm$5.27
319.09
$\pm$10.40
188.17
$\pm$11.23
446.02
$\pm$15.82
413.10
$\pm$10.78
228.82
$\pm$5.96
200.93
$\pm$7.89
374.51
$\pm$11.13
250.79
$\pm$8.63
Cora PubMed WikiCS WN18RR FB15K237 HIV PCBA Avg.
GCN
75.65
$\pm$1.37
75.61
$\pm$2.10
75.28
$\pm$1.34
73.79
$\pm$0.39
82.22
$\pm$0.28
64.84
$\pm$4.78
71.32
$\pm$0.49
73.76
GAT
76.24
$\pm$1.62
74.86
$\pm$1.87
76.28
$\pm$0.78
80.16
$\pm$0.27
88.93
$\pm$0.15
65.54
$\pm$6.93
70.12
$\pm$0.89
75.44
DGI
72.10
$\pm$0.34
73.13
$\pm$0.64
75.32
$\pm$0.95
75.75
$\pm$0.59
81.34
$\pm$0.15
59.62
$\pm$1.21
63.31
$\pm$0.89
71.22
BGRL
71.20
$\pm$0.30
75.29
$\pm$1.33
76.53
$\pm$0.69
75.44
$\pm$0.30
80.66
$\pm$0.29
63.95
$\pm$1.06
67.09
$\pm$1.00
72.67
GraphMAE
73.10
$\pm$0.40
74.32
$\pm$0.33
72.61
$\pm$0.39
78.99
$\pm$0.48
85.30
$\pm$0.16
61.04
$\pm$0.55
63.30
$\pm$0.78
73.07
GFT
78.35
$\pm$1.07
73.39
$\pm$1.68
79.13
$\pm$0.32
90.87
$\pm$0.25
89.89
$\pm$0.27
72.16
$\pm$1.69
72.74
$\pm$1.23
79.50
GFT + EMA
79.44
$\pm$0.89
75.82
$\pm$1.32
79.46
$\pm$0.58
90.58
$\pm$0.43
89.75
$\pm$0.19
72.39
$\pm$1.52
73.04
$\pm$1.01
80.06
GFT + AP
79.69
$\pm$1.07
74.03
$\pm$0.92
78.05
$\pm$0.58
89.56
$\pm$0.18
89.05
$\pm$0.18
71.86
$\pm$1.53
71.48
$\pm$0.99
79.10
GFT + Reset
80.07
$\pm$0.91
73.56
$\pm$0.85
79.73
$\pm$0.63
91.18
$\pm$0.43
88.09
$\pm$0.23
72.79
$\pm$1.65
71.95
$\pm$0.85
79.62
GFT + PT
78.57
$\pm$0.86
75.09
$\pm$1.21
79.39
$\pm$0.89
88.63
$\pm$0.15
88.45
$\pm$0.17
71.01
$\pm$1.74
73.73
$\pm$1.12
79.26
GFT + SimVQ
77.61
$\pm$0.73
76.59
$\pm$1.07
76.68
$\pm$0.76
82.72
$\pm$0.53
82.03
$\pm$0.35
66.57
$\pm$1.35
69.90
$\pm$0.91
76.01
GFT + RGVQ
80.85
$\pm$0.73
77.46
$\pm$0.94
80.10
$\pm$0.52
91.32
$\pm$0.26
90.45
$\pm$0.31
74.10
$\pm$1.49
75.68
$\pm$0.99
81.42
Table 6: Cross-domain and cross-task performance measured in Accuracy in the pre-training and fine-tuning setting. Bold highlights the best performance.
Cora PubMed Citeseer Photo Computer WikiCS Ratings Roman Questions
GCN
75.65
$\pm$1.37
75.61
$\pm$2.10
73.14
$\pm$0.67
96.10
$\pm$0.46
93.99
$\pm$0.12
75.28
$\pm$1.34
53.80
$\pm$0.60
91.27
$\pm$0.20
79.02
$\pm$0.60
GAT
76.24
$\pm$1.62
74.86
$\pm$1.87
72.22
$\pm$0.84
96.60
$\pm$0.33
94.09
$\pm$0.37
76.28
$\pm$0.78
55.54
$\pm$0.51
90.63
$\pm$0.14
77.95
$\pm$0.51
GraphGPS
82.84
$\pm$1.03
79.94
$\pm$0.26
72.73
$\pm$1.23
95.06
$\pm$0.13
91.19
$\pm$0.54
78.66
$\pm$0.49
53.10
$\pm$0.42
82.00
$\pm$0.61
71.73
$\pm$1.47
SGFormer
84.50
$\pm$0.80
80.30
$\pm$0.60
72.60
$\pm$0.20
95.10
$\pm$0.47
91.99
$\pm$0.76
73.46
$\pm$0.56
48.01
$\pm$0.49
79.10
$\pm$0.32
72.15
$\pm$1.31
Exphomer
82.77
$\pm$1.38
79.46
$\pm$0.35
71.63
$\pm$1.19
95.35
$\pm$0.22
91.47
$\pm$0.17
78.54
$\pm$0.49
53.51
$\pm$0.46
89.03
$\pm$0.37
GQT
86.44
$\pm$1.58
81.60
$\pm$1.35
73.14
$\pm$1.26
94.46
$\pm$0.68
92.13
$\pm$0.23
80.03
$\pm$0.19
54.04
$\pm$0.12
89.85
$\pm$0.73
76.52
$\pm$1.52
GQT + RGVQ
88.34
$\pm$1.32
86.54
$\pm$1.41
81.25
$\pm$1.01
97.66
$\pm$1.05
95.67
$\pm$0.36
83.58
$\pm$0.66
55.16
$\pm$0.29
90.98
$\pm$0.66
78.26
$\pm$1.07
Pre-computation of Contrastive Set. Before training, RGVQ constructs for each node sets of positive and negative samples, based on both structural and feature similarity. This step is performed once and reused during training. To implement this, neighbors are first extracted by scanning the adjacency matrix, which requires $O(|\mathcal{E}|)$ time. Then compute feature distances between each node and all others will take $O(|\mathcal{V}|^{2}d)$ time. However, in practice, we adopt a sampling strategy: for each node, we sample $M$ non-neighbor nodes (where $M$ is a small constant, e.g., 100) and compute their feature similarity. This limits the total cost of semantic similarity computation to $O(|\mathcal{V}|\cdot M\cdot d)$, which is linear in the number of nodes. After collecting both structurally and semantically similar candidates, we perform top-$k$ selection for each node to finalize its positive sample set, costing $O(|\mathcal{V}|\log k)$ time in total. Negative samples are the same from all nodes. Thus, the overall time complexity of the contrastive set construction process is $O(|\mathcal{E}|+|\mathcal{V}|\cdot M\cdot d+|\mathcal{V}|\log k)$, which is linear in the number of nodes and edges under fixed $M$ and $k$.
@@
E. Results with Deviations
We run each methods 5 times and provide the mean performance with standard deviations for codebook utilization, transferability, and serialization ability in Table 5, Table 6, and Table 7 respectively.

### v2

@@ B.1 Dataset
We use both homophilous and heterophilous graphs in our experiments. To implement empirical study and evaluate codebook utilization, we use various datasets, including Cora (Bojchevski & Günnemann, 2017), CiteSeer, PubMed (Namata et al., 2012), Amazon-Computer, Amazon-Photo (Shchur et al., 2018; McAuley et al., 2015), WikiCS (Mialon et al., 2021), Amazon-Ratings (Platonov et al., 2023), and Roman-Empire (Platonov et al., 2023). To assess transferability, we use cross-task and cross-domain datasets. Specifically, we use Cora, PubMed, and WikiCS for node classification; WN18RR (Shang et al., 2019) and FB15K237 Li et al. (2024b) for link prediction; and HIV (Hu et al., 2021) and PCBA (Chen et al., 2024) for graph classification. Finally, we evaluate serialization ability using the same datasets employed for codebook utilization. Detailed dataset statistics are summarized in Table 4.
B.2 Baseline
@@
We use different baselines for the empirical study and three parts of our main experiments.
@@
Empirical Study and Codebook Utilization. We primarily adopt codebook mitigation strategies originally developed in the vision and language domains, including EMA (Łańcucki et al., 2020), affine parameters (Huh et al., 2023), codebook reset (Zeghidour et al., 2021), and pretrained encoders (Zhao et al., 2024). We further include SimVQ (Zhu et al., 2024), which addresses the codebook collapse via one-single MLP layer over the latent basis vectors. Additionally, we compare with HQA-GAE (Zeng et al., 2025), a recent graph VQ model that applies a hierarchical VQ structure and an annealing strategy for codeword selection.
@@
Transferability. To evaluate the effectiveness of RGVQ in learning transferrable graph tokens, we integrate it into a graph foundation model, i.e., GFT, and compare it with vanilla Graph VQ and its variants with different mitigation strategies. Moreover, we include supervised GNNs, i.e., GCN, GAT, and GIN, and graph self-supervised methods, i.e., DGI (Veličković et al., 2018), BGRL (Thakoor et al., 2021), GraphMAE (Hou et al., 2022), and GAINT (Chien et al., 2021). The supervised GNNs are trained directly on each target dataset, while the self-supervised methods and all GFT variants are pretrained on the full set of datasets and then fine-tuned per target.
@@
Serialization. To further evaluate the effectiveness of RGVQ in serialization, we integrate it into the Graph Quantized Transformer (GQT) (Wang et al., 2025), where discrete tokens serve as the input sequence to a vanilla Transformer backbone. We follow the original sequence reconstruction method and compare the performance of RGVQ-enhanced GQT against the original GQT, supervised GNNs, and graph transformers, including GraphGPS (Rampášek et al., 2022), SGFormer (Wu et al., 2023), Exphomer (Shirzad et al., 2023), and NodeFormer (Wu et al., 2022).
B.3 Implementation Details
Empirical Study. We provide the hyperparameters and experimental setup used in the empirical study of codebook perplexity. We jointly train the single-head VQ model and the GAT encoder using the link prediction and feature reconstruction tasks, along with the commitment loss and vocabulary loss. The task weights are set to 0.01, 100, 0.1, and 0.9, respectively. We train the model for 100 epochs and report the highest perplexity during the training process for each method and a specific codebook size $K$. For all methods, we utilize the kmeans initialization and orthogonal regulation for the codebook, with a regularization weight of 0.1. The GNN consists of 4 layers with a hidden dimension of 256. AdamW is utilized as the optimizer with a learning rate of 1e-4 and a weight decay of 1e-5. For affine parameters, we use Euclidean distance and set the codebook decay to 0.9. For codebook reset, the threshold of deadcode is set to 10. For pretraining encoder, we pretrain the GNN encoder for 50 epochs before the joint training.
@@ B.3 Implementation Details
Codebook Utilization. The implementation of all collapse mitigation strategies is the same with the empirical study. For HQA-GAE, we use one-head VQ model and use the same hidden dimension and number of GNN layers as other methods, while all remaining hyperparameters follow the original paper. For RGVQ, we set the codebook size $K$ to 512, and sample 20 contrastive examples for both the positive and negative sets. The training weights for the link prediction, node feature reconstruction, contrastive regularization, commitment loss, and vocabulary loss are set to 0.01, 100, 1, 0.1, and 0.9, respectively. We set the temperature for the Gumbel-Softmax trick to 0.1.
@@
Transferability. We use RGVQ as a plugin within the pretraining pipeline of GFT. Specifically, we retain the link, node feature, and tree embedding reconstruction tasks and integrate RGVQ with respective weights of 100, 1, 0.01, and 10. For all other configurations, including downstream fine-tuning in GFT, we strictly follow the settings from the original paper (Wang et al., 2024).
@@
Serialization. We pretrain RGVQ using the same configuration as in the codebook utilization experiments for 200 epochs. We then load the pretrained checkpoints, including the codebook, GNN encoder, and VQ module weights. We follow the hyperparameter settings from the original GQT paper (Wang et al., 2025).
Appendix C Additional Experiments
Empirical Study. We additionally provide the quantization results on Cora and Citeseer datasets. The results shown in Figure 6 suggest that codebook collapse is a systematic problem in Cora and Citeseer datasets, even though the mitigation strategies are applied.
Ablation Study. We also provide the additional ablation study to further evaluate the contribution of each proposed module in RGVQ in this section. Here we use the normalized perplexity, defined as the ratio of utilized codebook entries to the total size of the codebook. First, we evaluate how codebook size affects the codebook utilization of RGVQ on more datasets. The results are shown in Figure 7(a). Across all datasets, RGVQ consistently maintains high normalized perplexity as the codebook size increases, showing strong robustness to the choice of codebook size. Notably, even with a large codebook size of 512, the model utilizes over 50% of the codebook capacity across all datasets. Second, we investigate how the Gumbel–Softmax temperature $\tau$ affects normalized perplexity on more datasets. A lower temperature yields a codebook assignment distribution that is closer to one-hot. As shown in Figure 7(b), unlike some prior work (Zeng et al., 2025) that rely on temperature annealing, a relatively low and fixed temperature is sufficient to address the non-differentiability of deterministic VQ on selected datasets. Finally, we examine how the number of contrastive samples in structure-aware regularization affects normalized perplexity. As shown in Figure 7(c), even a small number of contrastive samples (e.g., $n=10$ or $50$) achieves relatively high normalized perplexity, indicating effective codebook use. Increasing $n$ does not consistently lead to better utilization and may even degrade in some datasets, potentially due to added training noise.
In addition, we perform a comprehensive ablation study to evaluate the contribution of each component, with results reported in Table 5. We find that removing either the reparameterization or the structure-aware regularization results in a clear performance drop across all datasets, underscoring the necessity of these proposed modules. Furthermore, excluding either the structurally similar set or the feature-similar set reduces quantization diversity and consequently degrades downstream accuracy. These findings indicate that both topological and feature information play essential roles in determining the process of codebook selection.
(a) Codebook size vs. Perplexity.
(b) Temperature $(\tau)$ vs. Perplexity.
(c) Sample Number vs. Perplexity.
Figure 7: Ablation study results.
Table 5: Additional ablation study.
Feature Rec. Structure Rec. Reparam. Structure Sim. Feature Sim. Cora PubMed WikiCS
Perp. Acc. Perp. Acc. Perp. Acc.
94.47 78.35 4.14 73.39 10.18 79.13
172.32 79.87 215.35 76.32 153.35 79.84
135.45 79.12 208.16 76.29 179.49 79.79
211.69 80.85 319.09 77.46 228.82 80.10
Appendix D Complexity Analysis
@@
Pre-computation of Contrastive Set. Before training, RGVQ constructs for each node sets of positive and negative samples, based on both structural and feature similarity. This step is performed once and reused during training. To implement this, neighbors are first extracted by scanning the adjacency matrix, which requires $O(|\mathcal{E}|)$ time. Then compute feature distances between each node and all others will take $O(|\mathcal{V}|^{2}d)$ time. However, in practice, we adopt a sampling strategy: for each node, we sample $M$ non-neighbor nodes (where $M$ is a small constant, e.g., 100) and compute their feature similarity. This limits the total cost of semantic similarity computation to $O(|\mathcal{V}|\cdot M\cdot d)$, which is linear in the number of nodes. After collecting both structurally and semantically similar candidates, we perform top-$k$ selection for each node to finalize its positive sample set, costing $O(|\mathcal{V}|\log k)$ time in total. Negative pairs are sampled from the set of all nodes excluding the positives. Thus, the overall time complexity of the contrastive set construction process is $O(|\mathcal{E}|+|\mathcal{V}|\cdot M\cdot d+|\mathcal{V}|\log k)$, which is linear in the number of nodes and edges under fixed $M$ and $k$.